OpenStudy (anonymous):

After 75 games of the 2009 baseball season, the Texas Rangers had a record of 40 wins and 35 losses for a winning percentage of .533. You realize that this data fits a binomial model and want to use the formula P(x) = {n! / [x! (n-x)!]} * p^x * q^(n-x) to determine the probability of how many games they will win. What is the expected number of games the Rangers will win in the next 5 games? A. 0.467 B. 0.533 C. 2.33 D. 2.67

OpenStudy (anonymous):

@austinL would you know anything about this one?

OpenStudy (austinl):

Oh sweet lord yeebus, no clue. So sorry :(

OpenStudy (anonymous):

haha right! probs and stats is kinda confusing lol

OpenStudy (kropot72):

Do you know how to calculate the probability of winning 1, 2, 3, 4 and 5 games using the given formula? For example the probability of winning one game out of five games is \[P(one\ win)=\left(\begin{matrix}5 \\ 1\end{matrix}\right)0.533^{1}\times0.467^{4}=5\times0.533\times0.467^{4}=you\ can\ calculate\] Having found the probabilities of winning 1, 2, 3, 4 and 5 games, the next step is to set up a probability distribution table that will enable the expected number of games that will be won in the next five games to be calculated. let me know if you need more help.

OpenStudy (anonymous):

okay im getting A is that correct?

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@kropot72 u their>?

OpenStudy (kropot72):

What did you get for the five values of probability?

OpenStudy (anonymous):

0.1....

OpenStudy (kropot72):

P(one win) = 0.1267 What do you get for P(two wins), P(three wins), P(four wins) and P(five wins)?

OpenStudy (anonymous):

B, i got B for my answer now!

OpenStudy (kropot72):

Please answer my question regarding the values of probability. I need to be sure that you understand the method of solving.

OpenStudy (anonymous):

well what I did was when I got 0.1267 I rounded it up, that was all i did.

OpenStudy (kropot72):

The rounding needs to be to at least three decimal places. You need to do the calculations for four more values of probability. Then it will be possible to construct a probability distribution table and use it to calculate the answer.

OpenStudy (anonymous):

ya but it would still be B though right? im in a real big rush to get this questions done, i dont have a calculator, and the one online sux.

OpenStudy (kropot72):

So you are really saying that you are guessing the answer. I can help you with a link to a suitable online calculator and guide you in how to use it if you like.

OpenStudy (anonymous):

well, myalgera is a pretty good link I use. if thats what u were gona give me, but it still doesnt give me a deffinete answer

OpenStudy (kropot72):

There are five more calculations needed to arrive at the answer. the next one of them is \[P(two\ wins)=\left(\begin{matrix}5 \\ 2\end{matrix}\right)0.533^{2}\times0.467^{3}=10\times0.533^{2}\times0.467^{3}=you\ can\ calculate\]